### v1

Abstract
Recent advances in Large Language Models (LLMs) have catalyzed the development of Large Multimodal Models (LMMs). However, existing research primarily focuses on tuning language and image instructions, ignoring the critical pretraining phase where models learn to process textual and visual modalities jointly. In this paper, we propose a new pretraining paradigm for LMMs to enhance the visual comprehension capabilities of LLMs by introducing a novel cross-modal comprehension stage. Specifically, we design a dynamically learnable prompt token pool and employ the Hungarian algorithm to replace part of the original visual tokens with the most relevant prompt tokens. Then, we conceptualize visual tokens as analogous to a “foreign language” for the LLMs and propose a mixed attention mechanism with bidirectional visual attention and unidirectional textual attention to comprehensively enhance the understanding of visual tokens. Meanwhile, we integrate a detailed caption generation task, leveraging rich descriptions to further facilitate LLMs in understanding visual semantic information. After pretraining on 1.5 million publicly accessible data, we present a new foundation model called Croc. Experimental results demonstrate that Croc achieves new state-of-the-art performance on massive vision-language benchmarks. To support reproducibility and facilitate further research, we release the training code and pre-trained model weights at https://github.com/deepglint/Croc.
1 Introduction
The rapid expansion of mobile networks has accelerated the generation of vast data volumes, presenting unprecedented opportunities for the development and application of Large Language Models (LLMs) (An et al., 2024; Touvron et al., 2023; Bai et al., 2023). Despite their effectiveness, LLMs are primarily confined to processing textual inputs. To expand their multimodal perceptual capabilities, there is an increasing research focus on Large Multimodal Models (LMMs) (Yin et al., 2023; Jin et al., 2024; Yang et al., 2023c), which are designed to process and integrate inputs across multiple modalities.
As a milestone in LMM research, LLaVA (Liu et al., 2024b) leverages the language-only capabilities of GPT-4 (Achiam et al., 2023) to generate multimodal language-image instruction-following datasets, demonstrating impressive multimodal conversational capabilities. Building on this groundwork, LLaVA-1.5 (Liu et al., 2024a) enhances performance through simple modifications to the original LLaVA framework and incorporates academically oriented Visual Question Answering (VQA) datasets with structured response formatting prompts. In parallel, BLIP-2 (Li et al., 2023a) and MiniGPT-4 (Zhu et al., 2023) connect a frozen pre-trained vision encoder and a language model through a trainable Q-Former or a linear layer, effectively mapping image features into the input embedding space of the language model. Nonetheless, these methods achieve only a superficial integration of image features within the language model’s embedding space. In contrast, CogVLM (Wang et al., 2023) introduces a trainable visual expert module into the attention and Feed-Forward Network (FFN) layers of the language model. Despite this innovation, the freezing of the LLM limits its capability to attain an in-depth understanding of visual features directly.
Recent research highlights the crucial role of the pretraining process in LMMs. Flamingo (Alayrac et al., 2022) synergistically integrates pre-trained vision and language models through pretraining on comprehensive multimodal web corpora that combine text and image, enabling the performance of diverse multimodal tasks such as captioning, visual dialogue, and visual question-answering. However, the usage of large-scale pretraining data results in substantial resource consumption. VILA (Lin et al., 2024) reveals that the pretraining process substantively augments various model capabilities, including multi-image reasoning, improved in-context learning, and enriched world knowledge. Nevertheless, VILA uses 50M of the interleaved pre-training corpus to improve data diversity, which is more computationally expensive than LLaVA-1.5. LaVIT (Jin et al., 2023) introduces a meticulously designed visual tokenizer to transform non-linguistic images into a sequence of discrete tokens, thereby rendering them analogous to a foreign language that is interpretable by LLMs. However, this direct input of visual tokens into LLMs and fostering visual understanding through next-token prediction encounters significant hurdles. This limitation primarily originates from the inherent discrepancies between visual and textual tokens, particularly the absence of a robust causal linkage between sequential visual tokens. Furthermore, the application of unidirectional attention mechanisms in this context further restricts the LLM’s capacity to effectively comprehend discrete visual tokens.
In this paper, we introduce a novel pretraining paradigm for LMMs designed to significantly enhance the visual comprehension capabilities of LLMs by incorporating a pioneering cross-modal comprehension stage. Specifically, we design a dynamically learnable prompt token pool and apply the Hungarian algorithm to selectively replace part of the original visual tokens with the most relevant prompt tokens. Then we conceptualize visual tokens as analogous to a “foreign language” for the LLMs and propose a mixed attention mechanism with bidirectional visual attention and unidirectional textual attention to improve the understanding of visual tokens. Meanwhile, we integrate a detailed caption generation task, leveraging rich image descriptions to further facilitate LLMs in understanding visual semantic information. Experiment results demonstrate that our proposed Croc model achieves new state-of-the-art performance across multiple benchmarks. The main contributions of this paper are summarized as follows:
We introduce a new pretraining paradigm to enhance the visual comprehension capabilities of LLMs by introducing a novel cross-modal comprehension stage. This stage integrates visual token reconstruction and targets detailed caption generation.
For visual token reconstruction, we design a dynamically learnable prompt token pool and employ the Hungarian algorithm to replace part of the original image tokens with the most relevant prompt tokens. In addition, we propose a mixed attention mechanism with bidirectional visual attention and unidirectional textual attention for more comprehensive visual token understanding.
Experimental results demonstrate that our proposed Croc model achieves new state-of-the-art across various benchmarks and exhibits robust visual understanding and reasoning capabilities. We will release the training code and models to facilitate future research.
2 Related Work
Large Multimodal Model Pre-training. As a significant advancement, LLaVA (Liu et al., 2024b) meticulously filters the CC3M (Changpinyo et al., 2021) dataset down to 595K and maintains the frozen state of both the visual encoder and the LLM weights, exclusively training the projection layer to align features from the visual encoder and the LLM. However, this strategy predominantly results in limited deep feature integration between the visual encoder and the LLM, primarily due to the constraints imposed by the projection layer. To address this limitation, CogVLM (Wang et al., 2023) introduces a trainable visual expert module into the attention and feed-forward network layers of the language model. Despite this innovation, constrained by the frozen state of the LLM, it continues to face challenges in comprehending the “foreign language” of visual tokens. Recently, LaVIT (Jin et al., 2023) introduces a well-designed visual tokenizer to convert non-linguistic images into a sequence of discrete tokens. However, directly inputting visual tokens into LLM to enhance visual understanding of LLM through next-token prediction presents significant limitations. VILA (Lin et al., 2024) proposes an interleaved pertraining stage to augment the LLM to support visual input, but it relies on a 50M pertraining dataset, necessitating considerable computational resources.
Cross-Model Comprehension. To improve the performance of Masked Image Modeling (MIM) based pre-training methods, MVP (Wei et al., 2022) initially explores the integration of multimodal pre-training within the MIM framework. Subsequently, diverging from conventional methods that predominantly predict raw pixels or low-level features, MILAN (Hou et al., 2022) adopts an innovative approach by reconstructing image features infused with substantial semantic content derived from caption supervision. UnMasked Teacher (Li et al., 2023b) selectively masks video tokens exhibiting low semantic content and aligns the remaining unmasked tokens through a linear projection to their counterparts from the teacher model. Experimental results confirm that this approach achieves state-of-the-art performance across various video-related tasks. In a recent study, RILS (Yang et al., 2023b) introduces a novel pre-training framework that employs masked visual reconstruction within a language semantic space. This framework facilitates the extraction of structured information by vision models through the accurate semantic prediction of masked tokens. Meanwhile, EVA (Fang et al., 2023) demonstrates that recovering the masked-out tokenized semantic vision features is an efficient strategy for vision-centric representation learning, obviating the need for semantic feature quantization or further tokenization. Inspired by the above works, we propose a visual token reconstruction task to improve the visual comprehension capability of LLMs.
3 Methodology
3.1 Preliminaries of LLaVA and LLaVA-1.5
As the seminal work of visual instruction tuning, LLaVA (Liu et al., 2024b) presents the first attempt to use language-only GPT-4 (Achiam et al., 2023) to generate multimodal language-image instruction-following data. The framework of LLaVA comprises three essential components: a Visual Encoder for transforming input images into distinct visual embeddings, a Projector for mapping visual embeddings into the textual embedding space, and a Large Language Model for processing both visual and textual tokens and generating corresponding responses. LLaVA utilizes a two-stage instruction-tuning process. In the first stage, image-text pairs are converted to the single-turn conversation which requests the assistant to describe the image. Given the input visual tokens and textual tokens , both and are fed into LLM to produce a coherent response. The ground-truth prediction answer is represented by the original caption . For a sequence of length , the probability of generating contextually original caption = is calculated as follows:
In the first stage, both the visual encoder and LLM weights are frozen and only the projection layer is updated. In the second stage, LLaVA keeps the visual encoder weights frozen and updates both the pre-trained weights of the projection layer and the LLM. With simple modifications to LLaVA, LLaVA-1.5 (Liu et al., 2024a) integrates CLIP ViT-L/14@336px (Radford et al., 2021) with an MLP projection and incorporates academic task-oriented VQA data with response formatting prompts, resulting in better multimodal comprehension capability.
3.2 Cross-Modal Comprehension
In this section, we propose a novel cross-modal comprehension pre-training method. Fig. 2 illustrates the training pipeline of the proposed Croc model. Unlike LLaVA, Croc includes an additional pre-training phase between stages 1 and 2. We conceptualize image tokens as a foreign language of the LLM and design a dynamically learnable prompt token pool to replace part of the original visual tokens with the most relevant prompt tokens. To facilitate the image token reconstruction, we design a mixed mechanism of bidirectional visual attention and unidirectional textual attention. Meanwhile, we introduce a detailed caption generation task to further enhance the LLM’s understanding of visual tokens.
Prompt Visual Token Generation. Inspired by EVA (Fang et al., 2023), we use LLM to reconstruct the masked visual tokens conditioned on visible image tokens. Given an image , we first extract visual features using the image encoder . Following LLaVA, we select the features before and after the last Transformer layer, and the visual projector translates the visual features into visual tokens . After that, unlike the previous works (He et al., 2022; Li et al., 2023b), we introduce a learnable prompt token pool to replace part of visual tokens with the most relevant prompt tokens. To ensure full utilization of the prompt tokens in the token pool, we use the Hungarian algorithm (Kuhn, 1955) to associate each masked visual token with a corresponding prompt token. We denote the prompt token pool as , where and represent the number of prompt tokens and the feature dimension. Under the mask ratio of , we get the set of masked visual tokens awaiting replacement. We pad with into a set of size . To find a bipartite matching between and , we search for a permutation of elements with the lowest cost:
Mixed Attention Mechanism. Due to the inherent disparities between visual and textual tokens, the causal interactions between visual tokens are significantly weaker than those observed between textual tokens. Inspired by the UNILM (Dong et al., 2019), we introduce a mixed cross-modal attention mechanism. As illustrated in Fig. 3, to enhance pre-training effectiveness and facilitate a more comprehensive assimilation of visual tokens by the LLM, we employ bidirectional attention among visual tokens and unidirectional attention among textual tokens. Such an arrangement not only allows mutual attention among visual tokens but also allows each textual token to attend to all visual tokens. Therefore, this attention configuration significantly enhances the efficacy of LLM in understanding and learning from visual tokens.
Detailed Caption Data. To improve LLM’s understanding of visual tokens, we propose to use detailed caption data for pre-training. As shown in Fig. 4, compared to brief captions, the detailed captions contain a greater wealth of semantic information about image details, thus providing better guidance for LLM to learn visual tokens more effectively.
Training Objectives. To improve the large language model’s ability to learn visual tokens, we introduce two specific objectives: Visual Token Reconstruction (VTR) and Detailed Caption Generation (DCG). In the prompt visual token generation step, we randomly replace some of the original visual tokens with tokens from our prompt token pool, thus obtaining the prompt visual tokens . Then we concatenate the prompt visual tokens with the instruction text tokens and feed them into an LLM to generate a response of length , the probability of generating the contextual response = can be calculated:
After receiving the response , we split the first 576 tokens to compute the visual token reconstruction loss and the remaining tokens to compute the detailed caption generation loss .
The visual token reconstruction loss is calculated as follows:
where represents the index set of replaced visual tokens. Meanwhile, we maximize the likelihood of text tokens by employing the auto-regressive language modeling objective:
The overall training loss is the combination of and :
where is a loss weight used to balance the influence of different losses.
3.3 Training Pipeline
The overall training pipeline of our Croc model is shown in Fig. 2. Building upon LLaVA-1.5, the Croc undergoes a two-stage pre-training procedure followed by instruction tuning.
Stage 1: Cross-Modal Alignment. Following LLaVA-1.5, we first pretrain the projection layer with the identical 558K pretraining dataset used in LLaVA-1.5 to align image features to the LLM embedding space. During training, we keep both the visual encoder and the LLM weights frozen.
Stage 1.5: Cross-Modal Comprehension. Building on the cross-modal alignment stage, we introduce the cross-modal comprehension phase as a subsequent pre-training stage. To facilitate comprehensive learning of visual tokens by the LLM, we pretrain the projection layer as well as the LLM in this stage. We select 1.2M detailed image-text pairs from the ShareGPT4V (Chen et al., 2024b) dataset, as the detailed captions (as shown in Fig. 4) can enhance the visual token learning of the LLM. To prevent the degradation of the inherent capabilities of LLM, we also incorporate 300K pure text data. Please refer to the appendix for more details.
Stage 2: Instruction Tuning. Following LLaVA-1.5, we freeze the visual encoder weights and update both the pre-trained weights of the projection layer and LLM to improve its visual question answering capabilities. We employ the identical 665K instruction dataset used in LLaVA-1.5.
4 Experiments
| Method | LLM | Res. | Pretrain | Finetune | VQAv2 | GQA | VizWiz | SciQA-I | TextVQA |
| BLIP-2 | Vicuna-13B | 2242 | 129M | – | 65.0 | 41.0 | 19.6 | 61.0 | 42.5 |
| InstructBLIP | Vicuna-7B | 2242 | 129M | 1.2M | – | 49.2 | 34.5 | 60.5 | 50.1 |
| InstructBLIP | Vicuna-13B | 2242 | 129M | 1.2M | – | 49.5 | 33.4 | 63.1 | 50.7 |
| Shikra | Vicuna-13B | 2242 | 600K | 5.5M | 77.4∗ | – | – | – | – |
| IDEFICS-9B | LLaMA-7B | 2242 | 353M | 1M | 50.9 | 38.4 | 35.5 | – | 25.9 |
| IDEFICS-80B | LLaMA-65B | 2242 | 353M | 1M | 60.0 | 45.2 | 36.0 | – | 30.9 |
| Qwen-VL | Qwen-7B | 4482 | 1.4B† | 50M | 78.8∗ | 59.3∗ | 35.2 | 67.1 | 63.8 |
| Qwen-VL-Chat | Qwen-7B | 4482 | 1.4B∗ | 50M | 78.2∗ | 57.5∗ | 38.9 | 68.2 | |
| LLaVA-1.5-7B | Vicuna-7B | 3362 | 558K | 665K | 78.5∗ | 62.0∗ | 50.0 | 66.8 | 58.2 |
| LLaVA-1.5-13B | Vicuna-13B | 3362 | 558K | 665K | 80.0∗ | 63.3∗ | 61.3 | ||
| Croc-7B | Vicuna-7B | 3362 | 558K+1.5M | 665K | 64.2∗ | 50.0 | 70.1 | 60.4 | |
| Croc-13B | Vicuna-13B | 3362 | 558K+1.5M | 665K | 80.7∗ | 57.1 | 72.7 | 60.8 |
| Method | LLM | Res. | Pretrain | Finetune | POPE | MME | MMB | MMB-CN | SEED | LLaVA-W | MM-Vet |
| BLIP-2 | Vicuna-13B | 2242 | 129M | – | 85.3 | 1293.8 | – | – | 46.4 | 38.1 | 22.4 |
| InstructBLIP | Vicuna-7B | 2242 | 129M | 1.2M | 86.1 | – | 36.0 | 23.7 | 53.4 | 60.9 | 26.2 |
| InstructBLIP | Vicuna-13B | 2242 | 129M | 1.2M | 78.9 | 1212.8 | – | – | – | 58.2 | 25.6 |
| Shikra | Vicuna-13B | 2242 | 600K | 5.5M | – | – | 58.8 | – | – | – | – |
| IDEFICS-9B | LLaMA-7B | 2242 | 353M | 1M | 81.9 | – | 48.2 | 25.2 | – | – | – |
| IDEFICS-80B | LLaMA-65B | 2242 | 353M | 1M | 66.0 | – | 54.5 | 38.1 | – | – | – |
| Qwen-VL | Qwen-7B | 4482 | 1.4B | 50M | – | – | 38.2 | 7.4 | 56.3 | – | – |
| Qwen-VL-Chat | Qwen-7B | 4482 | 1.4B | 50M | – | 1487.5 | 60.6 | 56.7 | 58.2 | – | – |
| LLaVA-1.5-7B | Vicuna-7B | 3362 | 558K | 665K | 85.9 | 1510.7 | 64.3 | 58.3 | 58.6 | 63.4 | 30.5 |
| LLaVA-1.5-13B | Vicuna-13B | 3362 | 558K | 665K | 85.9 | 1531.3 | 67.7 | 63.6 | 61.6 | 70.7 | 35.4 |
| Croc-7B | Vicuna-7B | 3362 | 558K+1.5M | 665K | 86.9 | 1526.4 | 67.6 | 59.7 | 64.1 | 71.9 | 34.9 |
| Croc-13B | Vicuna-13B | 3362 | 558K+1.5M | 665K | 87.8 | 1591.4 | 69.9 | 62.9 | 64.2 | 74.7 | 36.2 |
| Method | LLM | Res. | Pretrain | Finetune | MMStar | DocVQA | AI2D | RealWorldQA |
| BLIP-2 | Vicuna-13B | 2242 | 129M | – | – | – | – | – |
| InstructBLIP | Vicuna-7B | 2242 | 129M | 1.2M | – | – | – | – |
| InstructBLIP | Vicuna-13B | 2242 | 129M | 1.2M | – | – | – | – |
| Shikra | Vicuna-13B | 2242 | 600K | 5.5M | – | – | – | – |
| IDEFICS-9B | LLaMA-7B | 2242 | 353M | 1M | 21.6 | – | 42.2 | 42.1 |
| IDEFICS-80B | LLaMA-65B | 2242 | 353M | 1M | 26.1 | – | 54.8 | 49 |
| Qwen-VL | Qwen-7B | 4482 | 1.4B | 50M | 32.5 | 65.1 | 57.7 | 37.8 |
| Qwen-VL-Chat | Qwen-7B | 4482 | 1.4B | 50M | 37.5 | 62.6 | 62.3 | 49.3 |
| LLaVA-1.5-7B | Vicuna-7B | 3362 | 558K | 665K | 30.3 | 28.1 | 55.5 | 54.8 |
| LLaVA-1.5-13B | Vicuna-13B | 3362 | 558K | 665K | 32.8 | 30.3 | 60.8 | 55.3 |
| Croc-7B | Vicuna-7B | 3362 | 558K+1.5M | 665K | 37.7 | 32.5 | 59.3 | 56.3 |
| Croc-13B | Vicuna-13B | 3362 | 558K+1.5M | 665K | 37.9 | 33.2 | 59.5 | 57.1 |
4.1 Implementation Details and Evaluation Benchmarks
Following LLaVA-1.5, we utilize CLIP ViT-L/14@336px (Radford et al., 2021) as the visual encoder and the Vicuna (Chiang et al., 2023) 7b/13B model as the LLM. During the cross-modal alignment stage, the learning rate for the projection layer is set to . In the cross-modal comprehension stage, the learning rates are adjusted to for both the LLM and projection layer and for the prompt token pool and prediction layer. The random mask ratio is set as and the size of the prompt token pool is set to . In the instruction tuning stage, we adopt the learning rate of for LLM and the projection layer. Following previous works (Yang et al., 2023a; An et al., 2023; Gu et al., 2024; Yang et al., 2024), we employ AdamW (Loshchilov & Hutter, 2019) as the optimizer, initialized with a weight decay of . The parameters and are set to and , respectively. We train Croc on NVIDIA A100 (80G) GPUs.
To prove the effectiveness of the Croc model, we evaluate our models across various benchmarks, including 1) General Visual Question Answering: GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017b), VizWiz (Gurari et al., 2018), LLaVA-Bench (In-the-Wild) (Liu et al., 2024b), MM-Vet (Yu et al., 2024), RealWorldQA; 2) OCR-Related Question Answering: TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021); 3) Illusion Benchmarks: POPE (Li et al., 2023c); 4) Comprehensive Reasoning Benchmarks: MMBench (Liu et al., 2024c), MME (Yin et al., 2023), MMStar (Chen et al., 2024a), MM-Vet (Yu et al., 2024); 5) Science Visual Question Answering: SciQA (Lu et al., 2022), AI2D (Kembhavi et al., 2016);
4.2 Quantitative Analysis
In Tab. 1, we present a detailed comparison between our model and previous state-of-the-art methods across various benchmarks tailored for academic tasks. Croc outperforms existing models on 4/5 benchmarks. Specifically, compared to LLaVA-1.5-7B, Croc-7B demonstrates performance improvements of 2.0%, 2.2%, 3.3%, and 2.2% on VQAv2, GQA, SciQA-I, and TextVQA. Similarly, Croc-13B shows improvements of 0.7%, 0.7%, 3.5%, and 1.1% over LLaVA-1.5-13B on VQAv2, GQA, VisWiz, and SciQA-I. Additionally, we compare Croc with baselines on instruction-following benchmarks. As shown in Tab. 2, Croc achieves superior performance on 6/7 datasets. Notably, compared to LLaVA-1.5-7B, Croc-7B exhibits significant performance gains of 1.0%, 15.7, 3.3%, 1.4%, 5.5%, 8.5%, and 4.4% on POPE, MME, MMB, MMB-CN, SEED, LLaVA-W, and MM-Vet. Furthermore, we evaluate Croc’s performance on the multimodal understanding benchmarks in Tab 3. Croc-7B obtains significant improvements of 7.4%, 4.4%, 3.8%, and 1.5% over LLaVA-1.5-7B on MMStar, DocVQA, AI2D, and RealWorldQA, respectively. These performance improvements are primarily attributed to the proposed novel cross-modal comprehension pretraining stage, which significantly enhances the LLM’s capability to integrate and understand visual tokens.
4.3 Qualitative Analysis
Attributable to the cross-modal comprehension stage, our proposed Croc model adeptly captures fine-grained semantic information from images, significantly enhancing its visual understanding and reasoning capabilities. In the first example shown in Fig. 5, unlike LLaVA-1.5, our model captures more detailed information from images, such as specific vehicle models like Audi and lower shooting angles. As illustrated in the second example of Fig. 5, our Croc model efficiently identifies and describes detailed semantic information, such as the “parachute” under the aircraft.
4.4 Ablation Study
| Pool Size | GQA | TextVQA | MMB | SEED | MMStar |
| 1 | 62.3 | 56.6 | 65.1 | 61.3 | 34.8 |
| 576 | 63.8 | 58.4 | 66.2 | 62.5 | 33.8 |
| 2048 | 64.2 | 60.4 | 67.6 | 64.1 | 37.7 |
| 4096 | 63.9 | 58.8 | 66.5 | 63.1 | 34.7 |
| Mask Ratio | GQA | TextVQA | MMB | SEED | MMStar |
| 0.00 | 63.1 | 45.9 | 66.1 | 62.0 | 33.3 |
| 0.50 | 63.8 | 59.2 | 67.6 | 62.6 | 34.6 |
| 0.75 | 64.2 | 60.4 | 67.6 | 64.1 | 37.7 |
| 0.90 | 63.0 | 58.2 | 65.6 | 62.2 | 34.3 |
| Method | GQA | TextVQA | MMB | SEED | MMStar |
| Nearest Neighbors | 63.1 | 58.9 | 66.5 | 62.5 | 35.0 |
| Hungarian Matching | 64.2 | 60.4 | 67.6 | 64.1 | 37.7 |
| Attention Mechianism | GQA | TextVQA | MMB | SEED | MMStar |
| Unidirectional Attention | 63.4 | 58.7 | 66.6 | 62.5 | 35.2 |
| Mixed Attention | 64.2 | 60.4 | 67.6 | 64.1 | 37.7 |
| Caption Type | GQA | TextVQA | MMB | SEED | MMStar |
| Brief Caption | 62.5 | 57.1 | 66.4 | 62.9 | 36.1 |
| Detailed Caption | 64.2 | 60.4 | 67.6 | 64.1 | 37.7 |
| Stage1 | Stage1.5 | GQA | TextVQA | MMB | SEED | MMStar |
| ✗ | ✓ | 63.2 | 57.7 | 64.7 | 62.1 | 33.1 |
| ✓ | ✓ | 64.2 | 60.4 | 67.6 | 64.1 | 37.7 |
| Method | Data | GQA | SciQA-I | TextVQA | POPE | MME | MMB | MMB-CN | SEED | MMStar | ||
| LLaVA-1.5-7B | ✗ | ✗ | ✗ | 62.0 | 66.8 | 58.2 | 85.9 | 1510.7 | 64.3 | 58.3 | 58.6 | 33.3 |
| LLaVA-1.5-7B⋇ | ✓ | ✗ | ✗ | 62.3 | 68.2 | 57.0 | 86.1 | 1450.7 | 62.0 | 53.9 | 60.9 | 33.0 |
| Croc-7B | ✓ | ✓ | ✗ | 62.5 | 68.8 | 56.1 | 87.1 | 1458.5 | 64.3 | 57.6 | 60.4 | 33.9 |
| Croc-7B | ✓ | ✗ | ✓ | 63.3 | 70.5 | 58.5 | 86.9 | 1513.8 | 65.9 | 58.0 | 62.6 | 33.6 |
| Croc-7B | ✓ | ✓ | ✓ | 64.2 | 70.1 | 60.4 | 86.9 | 1526.4 | 67.6 | 59.7 | 64.1 | 37.7 |
Prompt Token Pool Size. The prompt token pool is designed to replace part of the original visual tokens with the most relevant prompt tokens. As shown in Fig. 6(a), the larger the number of tokens in the pool, the easier it is to generate detail captions, resulting in a lower training loss (). The decrease in training loss is saturated when the pool size is 2,048. When the pool size is 1, 75% of the image tokens are directly dropped and the training is too difficult, resulting in poor performance in Tab. LABEL:table:pool_size. We observe that increasing the token pool from 1 to 2,048 tokens improves performance. However, expanding the pool further to 4,096 tokens degrades performance because the visual token reconstruction task is too easy to improve the detailed caption generation task.
Visual Token Mask Ratio. The mask ratio of visual tokens directly affects the difficulty of the visual token reconstruction task, and thus significantly affects the effectiveness of pretraining. In Table. LABEL:table:mask_ratio, we report the results of experiments with different mask ratios. Similar to the observations with MAE (He et al., 2022), we find that using a 75% masking ratio yields optimal results in several downstream benchmarks. Lower ratios, e.g., 0%, make the pretraining task too easy, while higher ratios, e.g., 90%, make the reconstruction task too difficult. Both extremes lead to reduced pretraining effectiveness.
Nearest Neighbors v.s. Hungarian Matching. To improve the utilization rate of the tokens in the prompt token pool, we utilize Hungarian Matching to replace 75% of visual tokens with the most relevant prompt tokens. In Fig. 6(b), we present a comparative analysis of token utilization using both Nearest Neighbors and Hungarian Matching. Due to the Hungarian Matching algorithm’s stringent requirement to select distinct prompt tokens, there has been a significant improvement in the overall utilization of prompt tokens. This enhancement of utilization substantially improves the representational capabilities of the model (Zhu et al., 2024). Therefore, as shown in Tab. LABEL:table:matching, employing the Hungarian Matching algorithm, in contrast to the Nearest Neighbors algorithm, yields substantial performance enhancements across all evaluated benchmarks.
Unidirectional Attention v.s. Mixed Attention. To verify the impact of the mixed attention mechanism, we conduct experiments to compare mixed attention with unidirectional attention used in LLaVA-1.5. As shown in Tab. LABEL:table:attention, our proposed mixed attention mechanism achieves significant performance improvement on all the benchmarks. To further explore the influence of different attention mechanisms, we compare the average attention scores of response tokens to visual tokens and language instruction tokens in Fig. 7. Here, the same question “Please describe in detail what is in the picture.” is used for 200 randomly selected images. The Croc-7B model exhibits significantly higher attention scores for visual tokens compared to the LLaVA-1.5-7B model, as indicated by the brighter region in the visual token part of the heatmap. This suggests that Croc-7B puts more emphasis on understanding and attending to the image content when generating responses.
Brief Caption v.s. Detailed Caption. To investigate the impact of detailed captions on our proposed cross-modal comprehension stage, we employ the BLIP model to generate brief captions for the 1.2M images from the ShareGPT4V dataset. As shown in Tab. LABEL:table:caption, the use of detailed captions achieves a significant performance improvement. This is mainly because the detailed caption contains rich semantic details of images (as shown in Fig. 4), facilitating visual token reconstruction by LLM.
Ablation on Stage, Data, and Objective. In Tab. LABEL:table:stage1, we present the performance results without the cross-modal alignment pretraining stage. The misalignment between the features of the LLM and the visual encoder significantly increases the difficulty of the visual token reconstruction task. Consequently, the performance of the model on all the benchmarks is reduced considerably.
In Tab. 5, we present a series of comprehensive ablation studies to elucidate the influence of pre-training data and different training objectives. Initially, we augment the 558K pertaining data of LLaVA-1.5 with the additional 1.5M data and apply the same training method as LLaVA-1.5. This augmentation leads to performance improvements in only 4/9 datasets, while the others exhibit declines. This phenomenon can be attributed to the significant alteration in the distribution of the original 558K pretraining dataset caused by the introduction of the additional data. Building upon this foundation, we incorporated a visual token reconstruction task, which facilitated the establishment of more effective integration between the LLM and visual tokens. This integration led to improved performance across 5/9 datasets. Additionally, benefiting from the detailed description captions, we observe performance improvements in 8/9 datasets after implementing the detailed caption generation task. Finally, by combining visual token reconstruction and detailed caption generation tasks, the detailed caption further enriches the LLM’s understanding of visual features, resulting in significant performance enhancements in our model.
| Method | Pretrain | Finetune | VQAv2 | GQA | VisWiz | SciQAI | POPE | MMB | MMB-CN | SEED | LLaVA-W | MM-Vet |
| VILA-7B | 50M | 1M | 79.9 | 62.3 | 57.8 | 68.2 | 85.5 | 68.9 | 61.7 | 61.1 | 69.7 | 34.9 |
| Croc-7B | 558K+1.5M | 802K | 80.1 | 63.5 | 55.2 | 72.3 | 86.9 | 69.1 | 60.5 | 63.0 | 73.3 | 36.8 |
| VILA-13B | 50M | 1M | 80.8 | 63.3 | 60.6 | 73.7 | 84.2 | 70.3 | 64.3 | 62.8 | 73.0 | 38.8 |
| Croc-13B | 558K+1.5M | 802K | 80.9 | 64.0 | 57.8 | 74.1 | 87.0 | 71.0 | 65.3 | 64.6 | 80.5 | 39.2 |
VILA v.s. Croc. As VILA incorporates an additional one million instruction data, we construct an 802K dataset for the instruction tuning stage to facilitate a more equitable comparison. Specifically, we expand the identical 665K instruction data used in LLaVA-1.5 to 802K by sampling an additional 142K instances from publicly accessible datasets. Please refer to the appendix for more details. As shown in Tab. 6, although our model only utilizes 1/25 of the pre-training data compared to VILA, Croc-7B achieves significant performance improvements in 8/10 of the benchmarks. Similarly, Croc-13B also achieves significant performance improvements in 9/10 of the benchmarks. These performance improvements demonstrate that our proposed cross-modal comprehension pretraining stage can facilitate the learning of visual tokens by LLMs, thereby substantially improving the visual comprehension and reasoning capacities of LMMs.
5 Conclusion
In this paper, we introduce a novel pretraining paradigm for LMMs to enhance the visual comprehension capabilities of LLMs. Our approach incorporates a new cross-modal comprehension stage designed to bridge the gap between the visual and textual domains. Specifically, we develop a dynamically learnable prompt token pool and apply the Hungarian algorithm to replace a portion of the original visual tokens with the most relevant prompt tokens. To further improve the model’s understanding, we propose a mixed attention mechanism that combines bidirectional visual attention with unidirectional textual attention, enabling a more comprehensive interpretation of visual tokens. Additionally, we incorporate a detailed caption generation task, utilizing rich descriptions to improve the LLM’s grasp of visual semantic information. Experimental results show that our method achieves state-of-the-art performance across multiple vision-language benchmarks. We hope that our work offers valuable insights for advancing large multimodal models.
Appendix A Appendix
A.1 Implementation Details
A.1.1 Hyperparameters
In Tab. 7, we present all the training hyperparameters used in different training stages. We use greedy decoding for evaluation to ensure reproducibility.
| Hyperparameter | Stage 1 | Stage 1.5 | Stage 2 |
| batch size | 256 | 128 | 128 |
| lr | 1e-3 | 1.5e-4/2e-5 | 2e-5 |
| lr schedule | cosine decay | cosine decay | cosine decay |
| lr warmup ratio | 0.03 | 0.03 | 0.03 |
| weight decay | 0 | 0 | 0 |
| epoch | 1 | 1 | 1 |
| optimizer | AdamW | AdamW | AdamW |
| DeepSpeed stage | 2 | 3 | 3 |
A.1.2 Datasets
In the cross-modal comprehension stage, to prevent the degradation of the inherent capabilities of LLM, we also incorporate 300K pure text data. Fig. 8(a) shows the composition of a dataset consisting of text-only data. This data includes 123K, 52K, 30K, 25K, 20K, 25K, 15K, and 10K samples from MathInstruct (Yue et al., 2023), Standford Alpaca (Taori et al., 2023), BELLE (BELLEGroup, 2023), OpenPlatypus (Lee et al., 2023), CodeAlpaca (Chaudhary, 2023), Firefly111https://github.com/yangjianxin1/Firefly, Webqa (Chang et al., 2022), Dolly (Mike et al., 2023).
To facilitate a more equitable comparison with VILA, we construct an 802K instruction dataset categorized into General, OCR, Chart, Math, Text-only, and Other. Fig. 8(b) illustrates the distribution of data across various categories. Specifically, the General dataset comprises 158K, 86K, 83K, 82.5K, 11K, 11K, and 9K samples sourced from LLaVA-Instruct (Liu et al., 2024b), VG (Krishna et al., 2017), VQAv2 (Goyal et al., 2017a), A-OKVQA (Schwenk et al., 2022), ShareGPT4V (Chen et al., 2024b), LAION-GPT4V 222https://huggingface.co/datasets/laion/gpt4v-dataset, and OKVQA (Marino et al., 2019). The OCR dataset includes 80K, 23K, 5K, 3K, and 2K samples from OCRVQA (Mishra et al., 2019), SynDog-EN (Kim et al., 2022), ST-VQA (Biten et al., 2019), IAM (Cheng et al., 2022), and InfoVQA (Mathew et al., 2022). Chart data is drawn from 10K, 18K, and 28K samples taken from DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2023), and DVQA (Kafle et al., 2018). To enhance the model logical capacity, we respective sample 5K data from Geomverse (Kazemi et al., 2023), Clevr (Johnson et al., 2017), IconQA (Lu et al., 2021b), RAVEN (Huang et al., 2023), Metamath (Yu et al., 2023) and append additional 3K, 1K, and 1K data from Goat (Liu & Low, 2023), Inter-GPS (Lu et al., 2021a), and LIMA (Zhou et al., 2024). Besides, we append 40K pure text from ShareGPT 333https://huggingface.co/datasets/RyokoAI/ShareGPT52K to prevent the degradation of the inherent capabilities of LLM and append 72K, 48K, and 2.5K other data from GQA (Hudson & Manning, 2019) and RefCOCO (Kazemzadeh et al., 2014), AI2D (Kembhavi et al., 2016).
A.2 More Qualitative Results
We present additional qualitative comparisons on visual understanding & reasoning (Fig. 9), image captioning (Fig. 10), and OCR (Fig. 11) tasks, respectively. Benefiting from our proposed cross-modal comprehension stage, the Croc model exhibits enhanced visual comprehension and reasoning capabilities. Meanwhile, it can accurately generate semantic image descriptions and demonstrate robust OCR capabilities.
A.3 Limitation Discussion
In this work, we have focused solely on image token reconstruction, which limits the scope to static images. However, for comprehensive video understanding, it is essential to consider both spatial and temporal token reconstruction. This would allow us to capture the dynamic changes that occur across frames and enhance the model’s ability to process and interpret video sequences more effectively. Expanding our approach to include spatial-temporal token reconstruction is a necessary step for future improvements in video analysis.
A.4 Ethic Discussion
Even though large multimodal models represent a major technological advancement with broad applications, their development and deployment bring significant ethical responsibilities. To ensure that the proposed Croc model benefits society while minimizing potential harm, it is essential to address key concerns such as bias, privacy, misinformation, environmental impact, intellectual property, accountability, and human oversight. These challenges are not unique to Croc but are pressing issues faced by the entire AI research and development community.